OpenStudy (anonymous):

A 13-foot ladder is leaning against a house when its base starts to slide away. By the time the base is 12 feet from the house, the base of the ladder is moving at a rate of 5ft/sec. At what rate is the area of the triangle formed by the ladder, wall, and ground changing at that time?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (kayders1997):

The 13 ft ladder is always going to stay the same in this problem and we know it's sliding down a wall at 5 ft/sec

OpenStudy (kayders1997):

You in calculus?

OpenStudy (anonymous):

Yeah, and I know that the 13 ft ladder is consistent and not changing. Just not sure of where to place all the values on the triangle (like what's DX/DT or DY/DT).

OpenStudy (anonymous):

First remember the area is \[ A=xy \]Now remember that \(z\) tells us the relationship between \(x\) and \(y\):\[ z^2=x^2+y^2 \]We already know \(z=13\), so: \[ y^2=13^2-x^2 \]

OpenStudy (anonymous):

The area is A = 1/2bh.

OpenStudy (anonymous):

We can differentiate with respect to time: \[ 2A=xy \implies 2A'=x'y+xy' \]

OpenStudy (anonymous):

Okay, change to \(2A=xy\) then.

OpenStudy (kayders1997):

sorry I have to think quick

OpenStudy (kayders1997):

You don't use area you use Pythagorean theorm

OpenStudy (anonymous):

We can get \(y'\) using our other equation:\[ y^2=13^2-x^2\implies 2yy' = -2xx' \implies y'= -\frac{xx'}{y} \]

OpenStudy (anonymous):

We can get \(y\) using: \[ y^2=13^2-x^2\implies |y|=y=\sqrt{13^2-x^2} \]

OpenStudy (anonymous):

They already gave us \(x\) and \(x'\), so we have enough information now to get \(y\), \(y'\), and finally \(A'\).

OpenStudy (kayders1997):

Ummm you don't have to make it this complicated you know x=12 12^2+y^2=13^2

OpenStudy (quickthing):

et tu, brute?

OpenStudy (kayders1997):

And thAn differentiate A=1/2(bh)

OpenStudy (anonymous):

Still confused...

OpenStudy (kayders1997):

Da/dt=1/2(dh/dt(b)+db/dt(h))

OpenStudy (kayders1997):

Base is db/dt so db/dt=-5

OpenStudy (kayders1997):

No positive 5

OpenStudy (kayders1997):

So so far we have da/dt=1/2(dh/dt(12)+5(h))

OpenStudy (kayders1997):

Did you find dh/dt already?

OpenStudy (anonymous):

No, but isn't h 13?

OpenStudy (kayders1997):

That's the hypotense were looking for the height opposite of the angle which would be side y, and that's not good that would start a whole new problem

OpenStudy (anonymous):

Then what's the purpose of 13?

OpenStudy (kayders1997):

We can find dh/dt quickly though we have to, to find the answer to this problem

OpenStudy (anonymous):

ok

OpenStudy (kayders1997):

So we know 13 stays all the same so we can plug that into the pythgegeon theorm x^2+y^2=z^2 so let's call z=13 so we can plug that in because it will always be the same x^2+y^2=13 now differentiate

OpenStudy (anonymous):

so y = 5?

OpenStudy (kayders1997):

Differentiate

OpenStudy (anonymous):

Oh, no. we have to find DY/DT

OpenStudy (anonymous):

ok

OpenStudy (kayders1997):

With respect to time

OpenStudy (kayders1997):

So you will get?

OpenStudy (kayders1997):

2x Dx/dt+?

OpenStudy (anonymous):

Yep! DY/DT = 2x(DX/DT)

OpenStudy (kayders1997):

Huh?

OpenStudy (anonymous):

2y(DY/DT) = 2x(DX/DT)?

OpenStudy (kayders1997):

Y^2 move the two out to the front subtract one from power 2y than differentiate with respect to time son 2ydy/dt so you have 2x Dx/dt+ 2y dy/dt =0 now put one of them on the other side

OpenStudy (kayders1997):

Oh yes sorry one just have to be -

OpenStudy (kayders1997):

Cause you moved it to the other side

OpenStudy (anonymous):

so -2xDX/DT?

OpenStudy (kayders1997):

Now we can do Pythagorean theorm to find y for the height

OpenStudy (kayders1997):

That works

OpenStudy (anonymous):

isn't y = 5?

OpenStudy (kayders1997):

Yes now we have all numbers somplug everything in

OpenStudy (kayders1997):

So Dx/dt = 5 y=5 x=12

OpenStudy (anonymous):

so -12?

OpenStudy (kayders1997):

That's what I got just don't forget units now we need to plug that in for the second part and we now know y=5

OpenStudy (anonymous):

so -119/2 ft2/sec?

OpenStudy (anonymous):

Thanks a bunch!

OpenStudy (kayders1997):

That's what I got :)

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

yep!

OpenStudy (kayders1997):

Good work @ducksonquack that was a really long process and you did a good job :)

OpenStudy (danjs):

The triangle lengths give \[y = \sqrt{13^2-x^2}\] Writing area in tersm of just the base \[A = x \sqrt{13^2-x^2}\] Differentiate w.r.t. time, dA/dt= ... you have the change in area in terms of just the base x and the rate of change of x now...

OpenStudy (kayders1997):

Yeah but it gives you the moment in time when the ladder is away from the house

OpenStudy (danjs):

They give you the instant to calculate, when x=12 and dx/dt=5

OpenStudy (kayders1997):

Right

OpenStudy (danjs):

That is most of these related rate problems from the book. rewrite in terms of single variable and differentiate the equation. Then use the info to calculate the instant in time they want.

OpenStudy (kayders1997):

I have the same question the teacher gave us the answer key just answers and these were both the answers